# PointDreamer: Zero-shot 3D Textured Mesh Reconstruction from Colored Point Cloud

### v2

@@ 3.4.3 Non-Border-First Unprojection (NBF)
The bottom of Figure 3 illustrates the idea of our proposed novel unprojection strategy namely Non-Border-First Unprojection. It begins with detecting border areas in multi-view images, so as to paint the texture atlas while deprioritizing them. After that, if some atlas pixels remain unpainted since their associated 3D points can not be seen from any non-border area, we then select from all areas, whether they are border areas or not, to paint the rest pixels. Now the question is, how to accurately detect border areas.
Interestingly, we find that, the borders between occluded and occluding parts of the shape in image $I_{k}$ (2D space) often correspond to the borders between visible and invisible areas in “view $k$ visible texture atlas $\mathcal{T}_{k}$” (UV space), where $\mathcal{T}_{k}$ is defined as a texture atlas in which only visible areas in view $k$ are filled and rest invisible areas are empty. See the bottom of Figure 3 as an example, where such border areas are marked by sky blue dashed lines. In posed images $I_{a}$ and $I_{b}$, these areas divide the occluded brown seat and the occluding gray bottom, while in view $a$ and $b$ visible texture atlas $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$, these areas divide the visible (painted) and invisible (transparent) parts in view $a$ and $b$, respectively. The reason for this correspondence of UV visibility border and 2D occlusion border is that, a continuous area within a chart of the texture atlas usually corresponds to a continuous segment of the 3D surface. When a chart contains both visible and invisible regions in a given view $k$, it often indicates that the corresponding 3D areas of invisible regions are occluded by some other part of the shape. Therefore, we design our Non-Border-First unprojection strategy as follows.
@@
    If some points cannot be seen from either view, we assign it with a view by direction priority considering all areas in all views whether they’re border areas or not.
@@
In this way, our non-border-first strategy can reasonably eliminate artifacts caused by inconsistent prediction in border areas. Altorithm 1 shows the corresponding pseudoCode. We conduct experiments to verify the effectiveness of our proposed strategy in Section 4.
@@ 4.3 Main Results
4.3.1 Qualitative Results
We present visual results in Figure 4. Since DHSP3D is much slower than other methods (4-6 hours per shape compared to $<$1 min), we randomly select two objects from GSO dataset for efficiency during comparison. We highly recommend referring to our supplementary materials for more visual results and a 360° video.
@@ 4.3.2 Quantitative comparisons
We summarize the quantitative results in Tables 4.2, and geometry evaluation results in our supplementary file. Our method outperforms baselines on most metrics and datasets, especially on the two perceptual metrics FID and LPIPS with a significant margin. However, regarding PSNR, which prioritizes pixel-level accuracy and thus may differ from human perception of quality, our method is sometimes outperformed by SPR. PSNR has been found to prefer blurry images [49]. We assume that SPR predicts a point’s color by blending nearby points’ colors, leading to blurry textures but avoiding extreme pixel-level errors. On the contrary, our method dreams the colors by its diffusion prior. While achieving visually better results, a few pixels with extremely high errors may lower the PSNR score.
4.4 Multi-view Consistency
Figure 4 (e) and our supplementary video show that, our method achieves high multi-view consistency. The reason is two-fold: (1) Unlike image-conditioned 3D generation, it’s easier to maintain multi-view consistency when reconstructing point clouds, since the input point cloud already determines the global shape and colors, leaving only local details to be dreamed by the model. (2) For inconsistencies in local details, our proposed Non-Border-First unprojection effectively addresses this issue; see Figure 7 (c) and Table 4.4.
@@ 4.4 Multi-view Consistency
Figure 7: Sub-Module Replacement Results: (a) Geometry module. (b) Inpainting module. (c) Unprojection module.
@@ 4.5.1 Anti-Noise Ability Analysis
We compare our method and baseline methods by applying them on point clouds with manually added Gaussian noise (standard deviation = 0.005) from the chair category of ShapeNetCoreV2 dataset. Table 4.4 shows the results, where our method, even with noisy input, outperforms baselines with clean input. Figure 5 (b) also shows our PointDreamer’s strong anti-noise ability. While baseline methods exhibit significant performance drops given noisy input, our PointDreamer’s performance degradation is minimal: even barely noticeable by human perception.
@@ 4.5.2 Sparsity Test
To evaluate our method’s ability to deal with sparse input, we decrease the input point number from 30k to 10k gradually, and present the results in Table 4.4. We can see only a small performance drop, and even with only 20k points as input, our method outperforms baseline methods with 30k points as input. When the input point number is decreased to 10k, our PointDreamer achieves higher PSNR and SSIM metrics compared to the Texture Field baseline with 30k points. We also provide visual comparisons in Figure 5 (c), where our method significantly outperforms baseline methods with only 10k points as input.
@@ 4.5.3 Texture Completion Ability
4.6 Method Component Analysis
We here compare different implementations for sub-modules (geometry extraction, inpainting, and unprojection module) of our method by replacing one module at a time while keeping others unchanged. All experiments in this subsection are conducted on the full test set of the chair category of ShapeNetCoreV2 dataset with Gaussian noise (standard deviation = 0.005).
4.6.1 Geometry Extraction
In addition to POCO, we further explore extracting geometry by SPR and depth inpainting to investigate how geometry quality influences the final reconstructed textured mesh. Please refer to our supplementary file for more details on depth inpainting and quantitative results. Figure 7 (a) shows that noisy geometry leads to noisy texture, and our adopted POCO performs the best.
4.6.2 2D Inpainting
We compare different inpainting modules in our pipeline, including nearest interpolation, linear interpolation, DiffPIR [53], and our adopted DDNM, where DiffPIR is another diffusion-based 2D image restoration method with inpainting ability. We provide the visual comparisons in Figure 7 (b) and quantitative results in our supplementary file. As can be seen, the other three methods, though perform overall reasonably, fail to produce as clear textures as DDNM, indicating the importance of a strong inpainting module to the reconstruction performance.
4.6.3 Unprojection
We present the comparison results of different unprojection strategies in Figure  7 (c) and Table 4.4.
  • Opt. Scratch: optimize the texture atlas from scratch;
@@ 4.6.3 Unprojection
  • Opt. Naive: further optimize Naive’s obtained atlas;
  • Opt. NBF: further optimizes NBF’s obtained atlas while considering only loss in non-border areas.
Figure 7 (c) shows that the naive unprojection yields the above-mentioned artifacts from border areas, which can be slightly reduced by further optimization. Only our NBF and Opt. NBF nearly eliminate such artifacts. Table 4.4 also shows that our proposed NBF and Opt. NBF outperform other unprojection strategies.
Our NBF strategy can also work for other methods [51, 37] that reconstruct mesh textures from multiview images. For example, InstantMesh [51] reconstructs a textured 3D mesh from a single image, by first generating multiview images, and then predicting the mesh by a large reconstruction model. As mentioned, predicting colors by a network trained with 3D data can result in blurry textures, as shown in Figure 8 (a). While unprojecting the multiview images to the mesh can achieve clearer textures, as shown in Figure 8 (b), artifacts often occur at border areas, such as the border of the eye as marked in red. Our NBF strategy effectively solved such artifacts, as shown in (c). More examples are provided in (d-g).
@@ Appendix A Geometry Evaluation Results
To compare the reconstructed geometry quality of PointDreamer and baseline methods while ignoring textures, we report geometry evaluation results in Table F. Same as our manuscript, we use ShapeNetCoreV2 [44] (Chair, Car, Motorbike categories), Google Scanned Objects [45] and Omniobject3d [46] datasets. We follow POCO [4] to use commonly-used metrics of Chamfer L1-distance ×100 (CD), normal consistency (NC), and F-Score with threshold value 1% (FS). See POCO for more details about these metrics. Note that, as mentioned in our manuscript, we use POCO as the geometry extraction module for both our PointDreamer and the Texture Field baseline.
@@ Appendix B More Visual Results
Figure 10 shows more visual comparisons of our PointDreamer against baseline methods. Figure 11 shows more visual results of our PointDreamer, including projected sparse multi-view images and inpainted dense multi-view images. We hope this can help the readers better understand our pipeline.
@@ C.1.2 Results
We present the visual comparisons in our manuscript and quantitative results in Table F, respectively. As can be seen, POCO, as a state-of-the-art deep-learning-based surface reconstruction approach, outperforms the other two methods by a significant margin, especially regarding FID. Both SPR and Depth Inpainting suffer from noisy geometry and thus low-quality textures. This indicates that a higher-performance geometry extraction module contributes to a more refined reconstructed textured mesh.
@@
We provide the quantitative results of replacing our 2D inpainting module from DDNM [24] to nearest interpolation, linear interpolation, and DiffPIR [53] in Table F. DiffPIR is another 2D-diffusion-based image restoration approach with inpainting ability. As can be seen, DDNM outperforms the other three methods thanks to its strong inpainting ability, which is consistent with the visual comparisons from Figure 4 (b) in our manuscript.
@@ D.1 Effect of Different K Values (Number of Viewpoints for Projection and Inpainting)
To investigate the effect of different numbers of viewpoints (denoted as $K$) for projection and inpainting, we conduct experiments on the motorbike category or ShapeNetCoreV2 dataset by using different $K$ values including 6, 8, and 20. Table F shows the distribution of cameras for each setting, together with the quantitative results. We can see that more views contribute to a slightly higher reconstruction quality. Visual comparisons in Figure 13 also show that, an insufficient number of views would produce artifacts in invisible or occluded areas, thus impacting the performance.

### v3

@@ 3.4.3 Non-Border-First Unprojection (NBF)
Building upon the above analysis, we propose the Non-Border-First (NBF) Unprojection strategy to eliminate artifacts arising from inconsistent predictions in border areas. The core principle is to prioritize assigning colors from non-border regions across views when filling the texture atlas.
Border Area Detection. To prioritize non-border areas, we first need to detect border areas in each view. Since the direct detection of occlusion borders within multiview 2D images is challenging, we propose to shift the detection process to UV space. This approach leverages a key correspondence we observe empirically: occlusion borders in image $I_{k}$ (2D space) often correspond to visibility borders in “view $k$ visible texture atlas $\mathcal{T}_{k}$” (UV space), where $\mathcal{T}_{k}$ is defined as a texture atlas in which only visible areas in view $k$ are filled and rest invisible areas are empty.
For example, Figure 6 (1) illustrates view $a$ and view $b$’s visible texture atlas $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$, respectively. As the navy blue dashed arrows show, the bottom side of the brown seat in $I_{a}$ and $I_{b}$ are mainly used to paint corresponding pixels in chart ① of $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$, and the bottom side of the gray cone is mainly used to paint chart ②. Zooming in, we can see in Figure 6 (2) that, the representive chart ① in both $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$ can be divided into a brown area (visible) and an empty area (invisible), respectively. We mark their border areas in sky blue dashed lines and denote such borders as “visibility border”. Note that the brown area corresponds to the bottom side of the brown seat of the chair, and the empty area corresponds to the parts that are occluded by the gray cone, as we marked by arrows with text “Correspond!” in sky blue in Figure 6. In other words, the invisible/visible areas in $\mathcal{T}_{k}$ correspond to the occluded/occluding areas in $I_{k}$. The reason for this correspondence is that, a continuous area within a chart of the texture atlas usually corresponds to a continuous segment of the 3D surface. When a chart contains both visible and invisible regions in a given view $k$, it often indicates that the corresponding 3D areas of invisible regions are occluded by some other part of the shape.
Border-Priority-Based Unprojection. Exploiting this correspondence, visibility borders in $\mathcal{T}_{k}$ (detected via edge detection algorithms) serve as proxies for occlusion borders in $I_{k}$. Consequently, NBF prioritizes non-border regions during texturing. For instance, in Figure 6 (3), example 3D point $\mathbf{p}$ and its corresponding pixel $I_{a}[i_{a},j_{a}]$ and $I_{b}[i_{b},j_{b}]$ in $I_{a}$ and $I_{b}$ are marked in small red circles. Since $I_{a}[i_{a},j_{a}]$ and $I_{b}[i_{b},j_{b}]$ are in border and non-border areas, respectively, the non-border-first principle guides us to pritize $I_{b}[i_{b},j_{b}]$ and paint point $\mathbf{p}$ as $I_{b}[i_{b},j_{b}]$’s color (brown). In this way, we effectively eliminate the gray artifact in naive or optimization unprojection caused by inconsistent prediction in border areas. Putting everything together, the detailed steps of the proposed NBF unprojection strategy are as follows.
@@
    If some points cannot be seen from either view, we assign each of them with a view by direction priority considering all areas in all views whether they’re border areas or not.
@@ 4.2 Main Results
4.2.1 Qualitative comparisons
We present visual results in Figure 7. Since DHSP3D is much slower than other methods (4-6 hours per shape compared to <<1 min), we randomly select two objects from GSO dataset for efficiency during comparison. We highly recommend referring to our supplementary materials for more visual results and a 360° video.
@@ 4.2.2 Quantitative comparisons
We summarize the quantitative results in Tables 4.1.3, and geometry evaluation results in our supplementary file. Our method outperforms baselines on most metrics and datasets, especially on the two perceptual metrics FID and LPIPS with a significant margin. However, regarding PSNR, which prioritizes pixel-level accuracy and thus may differ from human perception of quality, our method is sometimes outperformed by SPR. PSNR has been found to prefer blurry images [57]. We assume that SPR predicts a point’s color by blending nearby points’ colors, leading to blurry textures but avoiding extreme pixel-level errors. On the contrary, our method dreams the colors by its diffusion prior. While achieving visually better results, a few pixels with extremely high errors may lower the PSNR score.
4.2.3 Multi-view Consistency
Figure 7 (e) and our supplementary video show that, our method achieves high multi-view consistency. The reason is two-fold: (1) Unlike image-conditioned 3D generation, it’s easier to maintain multi-view consistency when reconstructing point clouds, since the input point cloud already determines the global shape and colors, leaving only local details to be dreamed by the model. (2) For inconsistencies in local details, our proposed Non-Border-First unprojection effectively addresses this issue; see Figure 11 (c) and Table 4.2.4.
4.2.4 More Visual Presentation
For readers to better understand our pipeline, we present the input point cloud, intermediate multiview images, and output meshes in Figure 8.
Figure 8: Visual presentation of the input point clouds, intermediate sparse and dense multiview images, and the output meshes.
@@ 4.2.4 More Visual Presentation
Figure 11: Ablation comparisons between our proposed NBF strategy and other unprojection strategies.
@@ 4.3.1 Anti-Noise Ability Analysis
We compare our method and baseline methods by applying them on point clouds with manually added Gaussian noise (standard deviation = 0.005) from the chair category of ShapeNetCoreV2 dataset. Table 4.2.4 shows the results, where our method, even with noisy input, outperforms baselines with clean input. Figure 9 (b) also shows our PointDreamer’s strong anti-noise ability. While baseline methods exhibit significant performance drops given noisy input, our PointDreamer’s performance degradation is minimal, even barely noticeable by human perception. We analyze that PointDreamer’s relative robustness to noisy input stems from the strong diffusion prior. Although the input sparse images are noisy, the adopted 2D diffusion model is trained on a vast amount of clean and plausible 2D images with no noise. Consequently, the model yields relatively clean inpainted images as it learns to generate samples that follow its training data’s distribution.
@@ 4.3.2 Sparsity Test
To evaluate our method’s ability to deal with sparse input, we decrease the input point number from 30k to 10k gradually, and present the results in Table 4.2.4. We can see only a small performance drop, and even with only 20k points as input, our method outperforms baseline methods with 30k points as input. When the input point number is decreased to 10k, our PointDreamer achieves higher PSNR and SSIM metrics compared to the Texture Field baseline with 30k points. We also provide visual comparisons in Figure 9 (c), where our method significantly outperforms baseline methods with only 10k points as input. We analyze that the high generation capability of the 2D diffusion prior is the main reason for this robustness to sparse input data.
@@ 4.3.3 Texture Completion Ability
4.4 Ablation Study on NBF Unprojection
To verify the effectiveness of our proposed NBF unprojection strategy, we conduct comparison experiments on the full test set of the chair category of ShapeNetCoreV2 dataset with Gaussian noise (standard deviation = 0.005). We present the comparison results in Figure 11 and Table 4.2.4, including the following methods:
  • Opt. Scratch: Randomly initialize a texture atlas and then optimize it by minimizing the per-pixel MSE loss between the mesh renderings and the inpainted multiview images, as described in Section 3.4.2;
@@ 4.4 Ablation Study on NBF Unprojection
  • Opt. Naive: Take Naive’s obtained atlas as initialization and optimize it by minimizing the per-pixel MSE loss between the mesh renderings and the inpainted multiview images, as described in Section 3.4.2;
  • Opt. NBF: Take NBF’s obtained atlas as initialization and optimize it by minimizing the per-pixel MSE loss between the mesh renderings and the inpainted multiview images’ non-border areas.
Figure 11 shows that the naive unprojection yields the above-mentioned artifacts from border areas, which can be slightly reduced by further optimization. Only our NBF and Opt. NBF nearly eliminate such artifacts. Table 4.2.4 also shows that our proposed NBF and Opt. NBF outperform other unprojection strategies.
Importantly, going beyond the specific colored-PC-to-mesh task, our proposed NBF unprojection strategy can effectively generalize to any method that textures an existing mesh given multiview images. In these methods, even with specifically designed techniques, the local border-area inconsistencies between multiview images and geometry can hardly be eliminated, which our NBF addresses.
We take InstantMesh [59] as an example. It reconstructs a textured 3D mesh from a single image, by first generating multiview images through a multiview diffusion model [61], and then outputting the mesh via a large reconstruction model that directly predict the mesh’s geometry and texture in 3D space. Note that the multiview diffusion technique is specifically designed for multiview-consistent generation.
We present visual comparisons in Figure 12. As mentioned in our introduction, learning colors in 3D space can result in blurry textures, as shown in Figure 12 (a). When unprojecting the multiview images to the mesh using the naive unprojection strategy, the texture becomes sharper, as shown in Figure 12 (b). However, as a cost, an artifact occurs in a border area, i.e. the border of the eye as marked in red. Our NBF strategy effectively solves this artifact, as shown in (c). More examples are provided in (d-g). These examples demonstrate that NBF can serve as a general solution to border-area artifacts in multiview-image-to-texture methods.
4.5 Real-life Data Experiments
Above, we followed the common routine of point cloud reconstruction works [47, 14, 10] to conduct experiments using point clouds sampled from existing meshes, because the calculation of existing quantitative evaluation metrics like PSNR require ground-truth meshes.
To better evaluate the effectiveness of our PointDreamer, we further capture point cloud inputs by scanning real-life objects with an Intel RealSense L515 LiDAR sensor. To ensure a more comprehensive evaluation, we select objects with diverse shapes and scales. As shown in Figure 13, the bottle in the top row is approximately a cylinder, the “Aoyu” in the second row and the box in the bottom row mainly consist of planes, and the bag in the third row exhibits a less regular shape. The bottle is the smallest in real life, leading to the noisiest input point cloud due to the limited sensor resolution. Since no ground-truth meshes are available for quantitative assessment of these scanned objects, we present the qualitative results in Figure 13.
Figure 13: Comparison results with real-scanned point clouds as input. Better zoom in.
As shown in Figure 13, other methods suffer from different levels of blurriness or jagged effects. Compared to them, our PointDreamer archives clearer texture reconstruction, thanks to the diffusion prior.
@@ Appendix A Geometry Evaluation Results
To compare the reconstructed geometry quality of PointDreamer and baseline methods while ignoring textures, we report geometry evaluation results in Table H. Same as our manuscript, we use ShapeNetCoreV2 [52] (Chair, Car, Motorbike categories), Google Scanned Objects [53] and Omniobject3d [54] datasets. We follow POCO [10] to use commonly-used metrics of Chamfer L1-distance ×100 (CD), normal consistency (NC), and F-Score with threshold value 1% (FS). See POCO for more details about these metrics. Note that, as mentioned in our manuscript, we use POCO as the geometry extraction module for both our PointDreamer and the Texture Field baseline.
@@ Appendix B More Visual Results
Figure 16 shows more visual comparisons of our PointDreamer against baseline methods.
@@ Appendix C Method Component Analysis
We here compare different implementations for our geometry extraction and inpainting sub-modules by replacing one module at a time while keeping everything else unchanged. All experiments in this section are conducted on the full test set of the chair category of ShapeNetCoreV2 dataset with Gaussian noise (standard deviation = 0.005).
@@ C.1.2 Results
We present the visual comparisons in Figure 14 (a) and quantitative results in Table H, respectively. As can be seen, POCO, as a state-of-the-art deep-learning-based surface reconstruction approach, outperforms the other two methods by a significant margin, especially regarding FID. Both SPR and Depth Inpainting suffer from noisy geometry and thus low-quality textures. This indicates that a higher-performance geometry extraction module contributes to a more refined reconstructed textured mesh.
@@
We compare different inpainting modules in our pipeline, including nearest interpolation, linear interpolation, DiffPIR [65], and our adopted DDNM, where DiffPIR is another diffusion-based 2D image restoration method with inpainting ability. We provide the visual comparisons in Figure 14 (b) and quantitative results in Table H. As can be seen, the other three methods, though perform overall reasonably, fail to produce as clear textures as DDNM, leading to a lower quantitative score. This indicates the importance of a strong inpainting module to the reconstruction performance.
@@ D.1 Effect of Different K Values (Number of Viewpoints for Projection and Inpainting)
To investigate the effect of different numbers of viewpoints (denoted as $K$) for projection and inpainting, we conduct experiments on the motorbike category or ShapeNetCoreV2 dataset by using different $K$ values including 6, 8, and 20. Table H shows the distribution of cameras for each setting, together with the quantitative results. We can see that more views contribute to a slightly higher reconstruction quality. Visual comparisons in Figure 18 also show that, an insufficient number of views would produce artifacts in invisible or occluded areas, thus impacting the performance.
@@ D.2.2 Sparsity Test
D.3 Comparison with recent image-to-3D, text-to-3D, and mesh texturing methods
D.3.1 Experimental Setting
We select two of the most updated and representative methods from each relevant category (mesh texturing, image-to-3D, and text-to-3D), which results in the following comparison baselines:
  • StableDreamFusion: A text-to-3D method based on SDS optimization. As the original DreamFusion is not open-source, we utilized an unofficial open-source implementation with substantial community support (8.6k GitHub stars).
  • ProlificDreamer: Another text-to-3D method leveraging SDS and further DMTet optimization.
  • Text-to-tex: A text-conditioned mesh texturing (text-to-texture) method that employs diffusion-based inpainting.
  • Easi-tex: An image-conditioned mesh texturing (image-to-texture) method, also utilizing diffusion-based inpainting.
  • CRM: An image-to-3D method built upon multiview diffusion and a large feed-forward reconstruction network.
  • StableFast3D (SF3D): Another image-to-3D approach, characterized utilizing large feed-forward reconstruction network.
The above methods involve three kinds of inputs:
  • Reference Image: For methods requiring a reference image, we rendered images from the ground-truth mesh of the object.
  • Untextured Mesh: For mesh texturing methods, we directly utilized the meshes reconstructed by our PointDreamer as the untextured mesh input.
  • Text Prompt: To ensure fair comparison for text-based methods, we initially generated prompts by feeding the rendered ground-truth images into a large language model (Qwen2.5-VL-32B-Instruct). These automatically generated prompts were then manually refined to provide more accurate and detailed descriptions of the objects. The complete text prompts are provided below:
    • Vintage analog clock: dark metal frame, matte khaki face with Arabic numerals. Hands at around 10:06:25. Small inset dial between center & numeral 12; ‘CROSLEY’ between center & numeral 6. Base with two legs, top ring for hanging.
    • A flat-sided wooden lion board with dual wooden wheels below. Features: yellow body, red petal mane, black eyes/orange nose/pink ears/blush marks. Green background & handle on lion’s back, ’Fisher-Price’ in white on red at base under the lion.
Note that most of these methods are relatively time-consuming (from about 15 minutes to more than an hour, compared to our method’s 100 seconds).
D.3.2 Experimental Results.
We present the comparison results in Figure 22, which shows:
  • The text-to-3D methods StableDreamFusion and ProlificDreamer generate significantly different results compared to the ground truth, since they only rely on text for generation without any visual information. Even when not considering fidelity, the generation results suffer from severe implausibility (e.g. ProlificDreamer’s generated clock) or even missing geometry (e.g. StableDreamFusion’s generated clock)
  • Text-to-tex is free from the geometry unplausibility issue with a mesh provided as input, but still it can hardly faithfully reconstruct the texture.
  • Easi-tex, though with an input image for reference, still struggles to generate high-fidelity textures, especially the backside without input information.
  • CRM generates higher-fidelity results then previous methods, but it still can not faithfully reconstruct the back view. In addition, its generated textures are blurry.
  • SF3D generates clearer textures from the front view, but its generated back views look highly implausible.
  • Our PointDreamer can reconstruct the objects with the highest fidelity, in other words, most similar to the ground truth object.
In summary, only our PointDreamer can achieve relatively high fidelity reconstruction, both theoretically and experimentally.